# TUTORIAL ARTICLE Contraction Theory for Nonlinear Stability Analysis and Learning-based Control: A Tutorial Overview

### v7

@@ 8.2 Parameter-Dependent Contraction Metric (aNCM)
where $B(x,\theta)=[b_{1}(x,\theta),\cdots,b_{m}(x,\theta)]$. We could define $\theta$ as $\theta^{\top},Z(\theta)^{\top}]^{\top}$ so we have $Y_{f}(q)\theta=f(q,\theta)$ and $Y_{b_{i}}(q)\theta=b_{i}(q;\theta)$. Such an over-parameterized system could be regularized using the Bregman divergence as in Boffi and Slotine [2020] (see Example 8.4), and thus we denote $\theta^{\top},Z(\theta)^{\top}]^{\top}$ as $\theta$ in the subsequent discussion.
@@
The adaptive Neural Contraction Metric (aNCM) in Fig. 13 is a DNN model for the optimal parameter-dependent contraction metric, given by solving the adaptive CV-STEM, i.e., (4.2) of Theorem 4.2 (or Theorem 4.6 for differential feedback) with its contraction constraint replaced by the following convex constraint:
@@
The aNCM given in Definition 8.1 has the following stability property along with its optimality due to the CV-STEM of Theorem 4.2 Tsukamoto et al. [2021].
@@
Suppose that Assumption 8.2 holds and let $\mathcal{M}$ define the aNCM of Definition 8.1. Suppose also that the true dynamics (8.1) is controlled by the following adaptive control law:
@@
where $d\mathcal{M}_{q}=[({\partial\mathcal{M}}/{\partial q_{1}})\mathtt{e},\cdots,({\partial\mathcal{M}}/{\partial q_{n}})\mathtt{e}]^{\top}/2$, $\mathtt{e}=x-x_{d}$, $\Gamma\succ 0$, $\sigma\in\mathbb{R}_{\geq 0}$, and $Y$, $_{d}, Y~\tilde{Y}$ are given in (8.20). Suppose further that the learning error in $\|\mathcal{M}-M\|\leq\epsilon_{\ell}$ of Theorem 6.3 additionally satisfies $\|d\mathcal{M}_{x_{d}}-dM_{x_{d}}\|\leq\epsilon_{\ell}$ and $\|d\mathcal{M}_{x}-dM_{x}\|\leq\epsilon_{\ell}$ in some compact set $\mathcal{S}$.
If $\exists\bar{b},\bar{\rho},\bar{y}\in\mathbb{R}_{>0}$ s.t. $\|B(x,\hat{\theta})\|\leq\bar{b}$, $\|R^{-1}(x,x_{d},u_{d})\|\leq\bar{\rho}$, $\|Y\|\leq\bar{y}$, $\|Y_{d}\|\leq\bar{y}$, and $\|\tilde{Y}\|\leq\bar{y}$ in (8.23) and (8.24), and if $\Gamma$ and $\sigma$ of (8.24) are selected to satisfy the following as in Theorem 8.1:
@@
for $\exists\alpha_{a}\in\mathbb{R}_{>0}$, $\alpha_{\ell}$ given in Theorem 6.3, and $\underline{m}$ and $\overline{m}$ given in $\underline{m}\mathrm{I}\preceq M\preceq\overline{m}\mathrm{I}$ of (2.32), then the system (8.1) is robust against bounded deterministic and stochastic disturbances, and we have the exponential bound (8.6) in the compact set $\mathcal{S}$. Furthermore, if $\epsilon_{\ell}=0$ (adaptive CV-STEM control), (8.24) with $\sigma=0$ guarantees asymptotic stability of $x$ to $x_{d}$ in (8.1).
@@
Replacing the contraction constraints of the CV-STEM in Theorem 4.2 by (8.21) and (8.22), the bound (8.6) and the asymptotic stability result can be derived as in the proof of Theorem 8.1 (see Theorem 4 and Corollary 2 of Tsukamoto et al. [2021] for details). Theorems 2.4 and 2.5 guarantee robustness of (8.1) against bounded deterministic and stochastic disturbances for $\sigma\neq 0$. $\qed$
@@
Let us again emphasize that, by using Theorem 4.6, the results of Theorems 8.1 and 8.2 can be extended to adaptive control with CCM-based differential feedback Lopez and Slotine [2021, 2021] (see Table 3 for the trade-offs).
Since the adaptation laws (8.4) and (8.24) in Theorems 8.1 and 8.2 yield an explicit bound on the steady-state error as in (8.6), it could be used as the objective function of the CV-STEM in Theorem 4.2, regarding $\Gamma$ and $\sigma$ as extra decision variables to get $M$ optimal in a sense different from Theorem 4.2. Smaller $\epsilon_{\ell}$ would lead to a weaker condition on them in (8.5) and (8.25). Also, the size of $\|\vartheta\|\leq\bar{\vartheta}$ in (8.6) can be adjusted simply by rescaling it (e.g., $\vartheta\to\theta/\bar{\vartheta}$). However, such a robustness guarantee comes with a drawback of having $\lim_{t\to\infty}\|\hat{\theta}(t)\|=0$ for $\sigma\neq 0$ in (8.4), leading to the trade-offs in different types of adaptation laws, some of which are given in Examples 8.28.4 (see also Remark 8.2).
@@
Let us briefly describe the following projection operator-based adaptation law, again for the Lagrangian system (8.13) of Example 8.1 with the unknown parameter vector $\theta$:
@@
See Lavretsky and Gibson [2012]. $\qed$
Since Lemma 8.1 guarantees the boundedness of $\|\hat{\theta}\|$, the adaptive controller (8.14), $u=-\mathcal{K}(\dot{\mathtt{q}}-\dot{\mathtt{q}}_{r})+Y\theta+Y\tilde{\theta}$, can be viewed as the exponentially stabilizing controller $\mathcal{K}(\dot{\mathtt{q}}-\dot{\mathtt{q}}_{r})+Y\theta$ (see Example 2.6) plus a bounded external disturbance $Y\tilde{\theta}$, implying robustness due to Theorem 2.4 Cai et al. [2006], Lavretsky and Gibson [2012], Chung et al. [2013], Lopez and Slotine [2021].
Let us also remark that, as in Example 8.1, the projection operator-based adaptation law (8.26) still achieves asymptotic stabilization. Applying the control law (8.14) with the adaptation (8.26) yields the following virtual system which has $q=\xi_{0}=[\dot{\mathtt{q}}_{r}^{\top},{\theta}^{\top}]^{\top}$ and $q=\xi_{1}=[\dot{\mathtt{q}}^{\top},\hat{\theta}^{\top}]^{\top}$ as its particular solutions:
@@
Using the convex property of the projection operator, i.e., $\tilde{\theta}^{\top}(\operatorname{Proj}{}(\hat{\theta},\xi,p)-\xi)\leq 0$ Cai et al. [2006], Lavretsky and Gibson [2012], this gives $\dot{V}_{s\ell}\leq-\int_{\xi_{0}}^{\xi_{1}}\delta q_{s}^{\top}\mathcal{K}\delta q_{s}$, which results in asymptotic stability of $\delta q$ due to Barbalat’s lemma [Slotine and Li, 1991, p. 405-406] as in Example 8.1.
@@
The dependence on $u$ and $\dot{\hat{\theta}}$ in $(d/dt)|_{\hat{\theta}}M$ can be removed by using $\partial_{b_{i}(x)}M+\partial_{b_{i}(x_{d})}M=0$ as in Theorem 8.1, and by using adaptation rate scaling introduced in Lopez and Slotine [2021]. In essence, the latter multiplies the adaptation law (8.24) by any strictly-increasing and strictly-positive scalar function $v(2\rho)$, and update $\rho$ as
@@
where $v_{\rho}=\partial v/\partial\rho$, $\eta\in\mathbb{R}_{>0}$, and $V_{\mathtt{e}}=\mathtt{e}^{\top}M(x,x_{d},u_{d},\hat{\theta})\mathtt{e}$ for $\mathtt{e}=x-x_{d}$ and $M$ given in Definition 8.1, so the additional term due to (8.29) cancels out the term involving $\dot{\hat{\theta}}$ in $(d/dt)|_{\hat{\theta}}\bar{W}$ of (8.21) (see Lopez and Slotine [2021] for details). Its robustness property follows from Theorem 8.2 also in this case.
@@
Using the Bregman divergence-based adaptation in Boffi and Slotine [2020], we could implicitly regularize the parameter estimate $\hat{\theta}$ as follows:
@@
where $d_{\psi}$ is the Bregman divergence defined as $d_{\psi}(x\|y)=\psi(x)-\psi(y)-(x-y)^{\top}\nabla\psi(y)$ for a convex function $\psi$, and $\mathcal{A}$ is a set containing only parameters that interpolate the dynamics along the entire trajectory. If $\hat{\theta}(0)=\mathrm{arg}\min_{b\in\mathbb{R}^{p}}\psi(b)$, we have $\lim_{t\to\infty}\hat{\theta}=\mathrm{arg}\min_{\vartheta\in\mathcal{A}}\psi(\vartheta)$, which regularizes $\hat{\theta}$ to converge to a parameter that minimizes $\psi$. For example, using $1$-norm for $\psi$ would impose sparsity on the steady-state parameter estimate $\hat{\theta}$ Boffi and Slotine [2020].
These extensions of adaptive control techniques described in Examples 8.28.4 could be utilized with contraction theory and learning-based control as in Theorems 8.1 and 8.2.
@@
Note that the results presented earlier in this section do not necessarily mean parameter convergence, $\lim_{t\to\infty}\|\tilde{\theta}\|=0$, as the adaptation objective is to drive the tracking error $\|x-x_{d}\|$ to zero [Slotine and Li, 1991, p. 331], not to find out the true parameter $\theta$ out of the many that achieve perfect tracking.
Asymptotic parameter convergence, $\lim_{t\to\infty}\|\tilde{\theta}\|=0$ for $\theta$ of (8.1) under Assumption 8.2, could be guaranteed if there is no disturbance and learning error with $\sigma=0$, and we have $\exists T,\alpha_{PE}\in\mathbb{R}_{>0}\text{ {\rm s.t.}{} }\int_{t}^{t+T}\tilde{Y}^{\top}\tilde{Y}d\tau\succeq\alpha_{PE}\mathrm{I},~\forall{t}$ for $\tilde{Y}$ given in (8.20) (the persistent excitation condition [Slotine and Li, 1991, p. 366]). We could also utilize the Bregman divergence-based adaptation to regularize the behavior of $\lim_{t\to\infty}\|\hat{\theta}\|$ as in Example 8.4.
9 Contraction Theory for Learned Models
@@
A typical approach is to perform system identification Shi et al. [2019], Singh et al. [2020], Manchester et al. [2021], Blocher et al. [2017] using trajectory data generated by (9.2):
@@
Let $q(0,t)=\xi_{0}(t)=x(t)$, $q(1,t)=\xi_{1}(t)=x^{*}(t)$, and $\textsl{g}=f_{L}$ in Theorems 5.2 and 5.3 as in Example 5.3, and define $\Delta_{L}$ as
@@
for the learning error condition $\|\Delta_{L}\|\leq\epsilon_{\ell 0}+\epsilon_{\ell 1}\|\xi_{1}-\xi_{0}\|$ in (5.3). If the function $f_{L}$ is learned to satisfy (5.3) with $\epsilon_{\ell 1}=0$, i.e., $\|\Delta_{L}\|\leq\epsilon_{\ell 0}$ in some compact set $\mathcal{S}$, and if there exists a contraction metric defined by $M$ bounded as $\underline{m}\mathrm{I}\preceq M\preceq\overline{m}\mathrm{I}$ as in (2.32), which renders (9.1) contracting as in (2.6) of Theorem 2.1 for deterministic systems, i.e.,
@@
and (2.46) of Theorem 2.5 for stochastic systems, i.e.,
@@
for $x$ and $x^{*}$ in (9.1) and (9.2), where $V_{\ell}=\int^{x}_{x^{*}}\|\Theta\delta x\|$ as in Theorem 2.3 with $M=\Theta^{\top}\Theta$. Furthermore, the systems (9.1) and (9.2) are robust against bounded deterministic and stochastic disturbances.
@@
Let $p_{t}^{*}=(x^{*},u(x^{*}(t),t),t)$ for notational simplicity. Since (9.2) can be written as $\dot{x}^{*}=f_{\mathrm{true}}(p_{t}^{*})=f_{L}(p_{t}^{*})+(f_{\mathrm{true}}(p_{t}^{*})-f_{L}(p_{t}^{*}))$ (see Example 5.3) and $\|\Delta_{L}\|=\|f_{\mathrm{true}}(p_{t}^{*})-f_{L}(p_{t}^{*})\|\leq\epsilon_{\ell 0}$, Theorem 2.4 holds with $\bar{d}=\epsilon_{\ell 0}$ as (9.1) is contracting, resulting in (9.6). Also, defining $\Delta_{L}$ as (9.3) in Theorems 5.2 and 5.3 results in robustness of (9.1) and (9.2) against bounded deterministic and stochastic disturbances due to (9.4) and (9.5), respectively. $\qed$
Theorem 9.1 is the theoretical foundation for stability analysis of model-free nonlinear dynamical systems. The bound (9.6) becomes tighter as we achieve smaller $\epsilon_{\ell 0}$ using more training data for verifying $\|\Delta_{L}\|\leq\epsilon_{\ell 0}$ (see Remark 5.1). From here onwards, we utilize contraction theory to provide stability guarantees to such model-free systems, partially enabling the use of the aforementioned model-based techniques.
9.1 Robust Control of Systems Modeled by DNNs
One challenge in applying Theorem 9.1 in practice is to find contraction metrics for the control non-affine nonlinear systems (9.1). This section delineates one way to construct a contraction metric in Theorem 9.1 for provably-stable feedback control, using the CV-STEM and NCM of Theorems 4.24.66.1, 6.36.5, and 7.1, along with the spectrally-normalized DNN of Definition 6.3.

### v8

@@ 8.2. Parameter-Dependent Contraction Metric (aNCM)
where $B(x,\theta)=[b_{1}(x,\theta),\cdots,b_{m}(x,\theta)]$. We could define $\theta$ as $\theta^{\top},Z(\theta)^{\top}]^{\top}$ so we have $Y_{f}(q)\theta=f(q,\theta)$ and $Y_{b_{i}}(q)\theta=b_{i}(q;\theta)$. Such an over-parameterized system could be regularized using the Bregman divergence as in boffi2020implicit (see Example LABEL:EX:ex:bregman), and thus we denote $\theta^{\top},Z(\theta)^{\top}]^{\top}$ as $\theta$ in the subsequent discussion.
@@
Adaptive Neural Contraction MetricsDef:aNCM The adaptive Neural Contraction Metric (aNCM) in Fig. 13 is a DNN model for the optimal parameter-dependent contraction metric, given by solving the adaptive CV-STEM, i.e., (4.2) of Theorem LABEL:THM:Thm:CV-STEM (or Theorem LABEL:THM:Thm:ccm_cvstem for differential feedback) with its contraction constraint replaced by the following convex constraint:
@@
The aNCM given in Definition LABEL:DEF:Def:aNCM has the following stability property along with its optimality due to the CV-STEM of Theorem LABEL:THM:Thm:CV-STEM ancm.
@@
Learning-based Adaptive Control with aNCMsThm:adaptiveNCMthm Suppose that Assumption 8.2 holds and let $\mathcal{M}$ define the aNCM of Definition LABEL:DEF:Def:aNCM. Suppose also that the true dynamics (8.1) is controlled by the following adaptive control law:
@@
where $d\mathcal{M}_{q}=[({\partial\mathcal{M}}/{\partial q_{1}})\mathtt{e},\cdots,({\partial\mathcal{M}}/{\partial q_{n}})\mathtt{e}]^{\top}/2$, $\mathtt{e}=x-x_{d}$, $\Gamma\succ 0$, $\sigma\in\mathbb{R}_{\geq 0}$, and $Y$, $_{d}, Y~\tilde{Y}$ are given in (8.20). Suppose further that the learning error in $\|\mathcal{M}-M\|\leq\epsilon_{\ell}$ of Theorem LABEL:THM:Thm:NCMstability2 additionally satisfies $\|d\mathcal{M}_{x_{d}}-dM_{x_{d}}\|\leq\epsilon_{\ell}$ and $\|d\mathcal{M}_{x}-dM_{x}\|\leq\epsilon_{\ell}$ in some compact set $\mathcal{S}$.
If $\exists\bar{b},\bar{\rho},\bar{y}\in\mathbb{R}_{>0}$ s.t. $\|B(x,\hat{\theta})\|\leq\bar{b}$, $\|R^{-1}(x,x_{d},u_{d})\|\leq\bar{\rho}$, $\|Y\|\leq\bar{y}$, $\|Y_{d}\|\leq\bar{y}$, and $\|\tilde{Y}\|\leq\bar{y}$ in (8.23) and (8.24), and if $\Gamma$ and $\sigma$ of (8.24) are selected to satisfy the following as in Theorem LABEL:THM:Thm:adaptive_robust_affine:
@@
for $\exists\alpha_{a}\in\mathbb{R}_{>0}$, $\alpha_{\ell}$ given in Theorem LABEL:THM:Thm:NCMstability2, and $\underline{m}$ and $\overline{m}$ given in $\underline{m}\mathrm{I}\preceq M\preceq\overline{m}\mathrm{I}$ of (2.32), then the system (8.1) is robust against bounded deterministic and stochastic disturbances, and we have the exponential bound (8.6) in the compact set $\mathcal{S}$. Furthermore, if $\epsilon_{\ell}=0$ (adaptive CV-STEM control), (8.24) with $\sigma=0$ guarantees asymptotic stability of $x$ to $x_{d}$ in (8.1).
@@
Replacing the contraction constraints of the CV-STEM in Theorem LABEL:THM:Thm:CV-STEM by (8.21) and (8.22), the bound (8.6) and the asymptotic stability result can be derived as in the proof of Theorem LABEL:THM:Thm:adaptive_robust_affine (see Theorem 4 and Corollary 2 of ancm for details). Theorems LABEL:THM:Thm:Robust_contraction_original and LABEL:THM:Thm:robuststochastic guarantee robustness of (8.1) against bounded deterministic and stochastic disturbances for $\sigma\neq 0$. $\qed$
@@
Let us again emphasize that, by using Theorem LABEL:THM:Thm:ccm_cvstem, the results of Theorems LABEL:THM:Thm:adaptive_robust_affine and LABEL:THM:Thm:adaptiveNCMthm can be extended to adaptive control with CCM-based differential feedback 9109296, lopez2021universal (see Table 3 for the trade-offs).
Since the adaptation laws (8.4) and (8.24) in Theorems LABEL:THM:Thm:adaptive_robust_affine and LABEL:THM:Thm:adaptiveNCMthm yield an explicit bound on the steady-state error as in (8.6), it could be used as the objective function of the CV-STEM in Theorem LABEL:THM:Thm:CV-STEM, regarding $\Gamma$ and $\sigma$ as extra decision variables to get $M$ optimal in a sense different from Theorem LABEL:THM:Thm:CV-STEM. Smaller $\epsilon_{\ell}$ would lead to a weaker condition on them in (8.5) and (8.25). Also, the size of $\|\vartheta\|\leq\bar{\vartheta}$ in (8.6) can be adjusted simply by rescaling it (e.g., $\vartheta\to\theta/\bar{\vartheta}$). However, such a robustness guarantee comes with a drawback of having $\lim_{t\to\infty}\|\hat{\theta}(t)\|=0$ for $\sigma\neq 0$ in (8.4), leading to the trade-offs in different types of adaptation laws, some of which are given in Examples LABEL:EX:ex:projection_operatorLABEL:EX:ex:bregman (see also Remark 8.2).
@@
Projection Operatorsex:projection_operator Let us briefly describe the following projection operator-based adaptation law, again for the Lagrangian system (8.13) of Example LABEL:EX:ex:lag_adaptive0 with the unknown parameter vector $\theta$:
@@
See lavretsky2012projection. $\qed$
Since Lemma 8.1 guarantees the boundedness of $\|\hat{\theta}\|$, the adaptive controller (8.14), $u=-\mathcal{K}(\dot{\mathtt{q}}-\dot{\mathtt{q}}_{r})+Y\theta+Y\tilde{\theta}$, can be viewed as the exponentially stabilizing controller $\mathcal{K}(\dot{\mathtt{q}}-\dot{\mathtt{q}}_{r})+Y\theta$ (see Example LABEL:EX:ex:lag_metric) plus a bounded external disturbance $Y\tilde{\theta}$, implying robustness due to Theorem LABEL:THM:Thm:Robust_contraction_original 1576865, lavretsky2012projection, Ref:phasesync, lopez2021universal.
Let us also remark that, as in Example LABEL:EX:ex:lag_adaptive0, the projection operator-based adaptation law (8.26) still achieves asymptotic stabilization. Applying the control law (8.14) with the adaptation (8.26) yields the following virtual system which has $q=\xi_{0}=[\dot{\mathtt{q}}_{r}^{\top},{\theta}^{\top}]^{\top}$ and $q=\xi_{1}=[\dot{\mathtt{q}}^{\top},\hat{\theta}^{\top}]^{\top}$ as its particular solutions:
@@
Using the convex property of the projection operator, i.e., $\tilde{\theta}^{\top}(\operatorname{Proj}{}(\hat{\theta},\xi,p)-\xi)\leq 0$ 1576865, lavretsky2012projection, this gives $\dot{V}_{s\ell}\leq-\int_{\xi_{0}}^{\xi_{1}}\delta q_{s}^{\top}\mathcal{K}\delta q_{s}$, which results in asymptotic stability of $\delta q$ due to Barbalat’s lemma [Ref_Slotine, p. 405-406] as in Example LABEL:EX:ex:lag_adaptive0.
@@
Adaptation Rate Scalingex:adaptation_rate_scale The dependence on $u$ and $\dot{\hat{\theta}}$ in $(d/dt)|_{\hat{\theta}}M$ can be removed by using $\partial_{b_{i}(x)}M+\partial_{b_{i}(x_{d})}M=0$ as in Theorem LABEL:THM:Thm:adaptive_robust_affine, and by using adaptation rate scaling introduced in lopez2021universal. In essence, the latter multiplies the adaptation law (8.24) by any strictly-increasing and strictly-positive scalar function $v(2\rho)$, and update $\rho$ as
@@
where $v_{\rho}=\partial v/\partial\rho$, $\eta\in\mathbb{R}_{>0}$, and $V_{\mathtt{e}}=\mathtt{e}^{\top}M(x,x_{d},u_{d},\hat{\theta})\mathtt{e}$ for $\mathtt{e}=x-x_{d}$ and $M$ given in Definition LABEL:DEF:Def:aNCM, so the additional term due to (8.29) cancels out the term involving $\dot{\hat{\theta}}$ in $(d/dt)|_{\hat{\theta}}\bar{W}$ of (8.21) (see lopez2021universal for details). Its robustness property follows from Theorem LABEL:THM:Thm:adaptiveNCMthm also in this case.
@@
Bregman Divergenceex:bregman Using the Bregman divergence-based adaptation in boffi2020implicit, we could implicitly regularize the parameter estimate $\hat{\theta}$ as follows:
@@
where $d_{\psi}$ is the Bregman divergence defined as $d_{\psi}(x\|y)=\psi(x)-\psi(y)-(x-y)^{\top}\nabla\psi(y)$ for a convex function $\psi$, and $\mathcal{A}$ is a set containing only parameters that interpolate the dynamics along the entire trajectory. If $\hat{\theta}(0)=\mathrm{arg}\min_{b\in\mathbb{R}^{p}}\psi(b)$, we have $\lim_{t\to\infty}\hat{\theta}=\mathrm{arg}\min_{\vartheta\in\mathcal{A}}\psi(\vartheta)$, which regularizes $\hat{\theta}$ to converge to a parameter that minimizes $\psi$. For example, using $1$-norm for $\psi$ would impose sparsity on the steady-state parameter estimate $\hat{\theta}$ boffi2020implicit.
These extensions of adaptive control techniques described in Examples LABEL:EX:ex:projection_operatorLABEL:EX:ex:bregman could be utilized with contraction theory and learning-based control as in Theorems LABEL:THM:Thm:adaptive_robust_affine and LABEL:THM:Thm:adaptiveNCMthm.
@@
Note that the results presented earlier in this section do not necessarily mean parameter convergence, $\lim_{t\to\infty}\|\tilde{\theta}\|=0$, as the adaptation objective is to drive the tracking error $\|x-x_{d}\|$ to zero [Ref_Slotine, p. 331], not to find out the true parameter $\theta$ out of the many that achieve perfect tracking.
Asymptotic parameter convergence, $\lim_{t\to\infty}\|\tilde{\theta}\|=0$ for $\theta$ of (8.1) under Assumption 8.2, could be guaranteed if there is no disturbance and learning error with $\sigma=0$, and we have $\exists T,\alpha_{PE}\in\mathbb{R}_{>0}\text{ {\rm s.t.}{} }\int_{t}^{t+T}\tilde{Y}^{\top}\tilde{Y}d\tau\succeq\alpha_{PE}\mathrm{I},~\forall{t}$ for $\tilde{Y}$ given in (8.20) (the persistent excitation condition [Ref_Slotine, p. 366]). We could also utilize the Bregman divergence-based adaptation to regularize the behavior of $\lim_{t\to\infty}\|\hat{\theta}\|$ as in Example LABEL:EX:ex:bregman.
9. Contraction Theory for Learned Models
@@
A typical approach is to perform system identification neurallander, 47710, cdc_systemid, 7992901 using trajectory data generated by (9.2):
@@
Contraction Theory for Learned ModelsThm:NCMstability_modelfree Let $q(0,t)=\xi_{0}(t)=x(t)$, $q(1,t)=\xi_{1}(t)=x^{*}(t)$, and $\textsl{g}=f_{L}$ in Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto as in Example LABEL:EX:ex:learning_prob3, and define $\Delta_{L}$ as
@@
for the learning error condition $\|\Delta_{L}\|\leq\epsilon_{\ell 0}+\epsilon_{\ell 1}\|\xi_{1}-\xi_{0}\|$ in (5.3). If the function $f_{L}$ is learned to satisfy (5.3) with $\epsilon_{\ell 1}=0$, i.e., $\|\Delta_{L}\|\leq\epsilon_{\ell 0}$ in some compact set $\mathcal{S}$, and if there exists a contraction metric defined by $M$ bounded as $\underline{m}\mathrm{I}\preceq M\preceq\overline{m}\mathrm{I}$ as in (2.32), which renders (9.1) contracting as in (2.6) of Theorem LABEL:THM:Thm:contraction for deterministic systems, i.e.,
@@
and (2.46) of Theorem LABEL:THM:Thm:robuststochastic for stochastic systems, i.e.,
@@
for $x$ and $x^{*}$ in (9.1) and (9.2), where $V_{\ell}=\int^{x}_{x^{*}}\|\Theta\delta x\|$ as in Theorem LABEL:THM:Thm:path_integral with $M=\Theta^{\top}\Theta$. Furthermore, the systems (9.1) and (9.2) are robust against bounded deterministic and stochastic disturbances.
@@
Let $p_{t}^{*}=(x^{*},u(x^{*}(t),t),t)$ for notational simplicity. Since (9.2) can be written as $\dot{x}^{*}=f_{\mathrm{true}}(p_{t}^{*})=f_{L}(p_{t}^{*})+(f_{\mathrm{true}}(p_{t}^{*})-f_{L}(p_{t}^{*}))$ (see Example LABEL:EX:ex:learning_prob3) and $\|\Delta_{L}\|=\|f_{\mathrm{true}}(p_{t}^{*})-f_{L}(p_{t}^{*})\|\leq\epsilon_{\ell 0}$, Theorem LABEL:THM:Thm:Robust_contraction_original holds with $\bar{d}=\epsilon_{\ell 0}$ as (9.1) is contracting, resulting in (9.6). Also, defining $\Delta_{L}$ as (9.3) in Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto results in robustness of (9.1) and (9.2) against bounded deterministic and stochastic disturbances due to (9.4) and (9.5), respectively. $\qed$
Theorem LABEL:THM:Thm:NCMstability_modelfree is the theoretical foundation for stability analysis of model-free nonlinear dynamical systems. The bound (9.6) becomes tighter as we achieve smaller $\epsilon_{\ell 0}$ using more training data for verifying $\|\Delta_{L}\|\leq\epsilon_{\ell 0}$ (see Remark 5.1). From here onwards, we utilize contraction theory to provide stability guarantees to such model-free systems, partially enabling the use of the aforementioned model-based techniques.
9.1. Robust Control of Systems Modeled by DNNs
One challenge in applying Theorem LABEL:THM:Thm:NCMstability_modelfree in practice is to find contraction metrics for the control non-affine nonlinear systems (9.1). This section delineates one way to construct a contraction metric in Theorem LABEL:THM:Thm:NCMstability_modelfree for provably-stable feedback control, using the CV-STEM and NCM of Theorems LABEL:THM:Thm:CV-STEMLABEL:THM:Thm:ccm_cvstemLABEL:THM:Thm:NCMstability1, LABEL:THM:Thm:NCMstability2LABEL:THM:Thm:ncm_ccm_clf, and LABEL:THM:Thm:lagros_stability, along with the spectrally-normalized DNN of Definition LABEL:DEF:Def:SN.